OpenStudy (anonymous):

what is the solution to this equation 3/8x+1= 9/4

OpenStudy (texaschic101):

Is it (3/8)x + 1 = 9/4 OR 3/(8x + 1) = 9/4

OpenStudy (anonymous):

the first one

OpenStudy (texaschic101):

3/8x + 1 = 9/4 -- if you want to make this easier, multiply everything by 8 3x + 8 = 18 3x = 18 - 8 3x = 10 x = 10/3 or 3/8x + 1 = 9/4 -- subtract 1 from both sides 3/8x = 9/4 - 1 3/8x = 9/4 - 4/4 3/8x = 5/4 x = (5/4) / (3/8) x = 5/4 * 8/3 x = 40/12 x = 10/3 same answer...just depends on how you want to do it

OpenStudy (anonymous):

thankss it helped alot

OpenStudy (texaschic101):

anytime :)